Fama's Llamas has a weighted average cost of capital of 12 percent. The company's cost of equity is 16.5 percent, and its pretax cost of debt is 8 percent. The tax rate is 31 percent. What is the company's target debt-equity ratio?
Answer #1

After-tax cost of debt=8*(1-tax rate)

=8*(1-0.31)=5.52%

Let debt be $x

Equity be $y

Total=$(x+y)

WACC=Respective costs*Respective weight

12=(x/(x+y)*5.52)+(y/(x+y)*16.5)

12*(x+y)=5.52x+16.5y

12x+12y=5.52x+16.5y

x=(16.5-12)y/(12-5.52)

=0.6944y

Hence debt-equity ratio=Debt/equity

=0.6944(Approx).